lilmami:

Given the linear equation 2x + y = 6, perform the necessary operations to put the equation into the proper general form. Explain in complete sentences how you knew that the equation was in the proper general form. Complete your work in the space provided or upload a file that can display math symbols if your work requires it. Include the entire process for establishing the general form of the equation and the general form.

dizzymustang156:

i can help

lilmami:

plz

dizzymustang156:

Step-by-step explanation: The general equation is y = mx + c... The given equation is 2x + y = 6 Firstly, move everything on the left except the y to the right.i.e. y = 6 - 2x Secondly, rearrange the values on the right to look just like the general equation. y = -2x + 6 The equation is now in proper general form because it matches the format laid down.

dizzymustang156:

there you go

lilmami:

thxxxxxxx

jhonyy9:

do you know the standard equation form of a line ?

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @dizzymustang156 Step-by-step explanation: The general equation is y = mx + c... The given equation is 2x + y = 6 Firstly, move everything on the left except the y to the right.i.e. y = 6 - 2x Secondly, rearrange the values on the right to look just like the general equation. y = -2x + 6 The equation is now in proper general form because it matches the format laid down. \(\color{#0cbb34}{\text{End of Quote}}\) good job

dizzymustang156:

thanx